OpenStudy (anonymous):

Can someone help me? I'm having a hard time with Algebra 1.. Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

are you supposed to come up with some hokey "word problem" for this, or just say it in english?

OpenStudy (anonymous):

I guess just say it in English haha :)

OpenStudy (anonymous):

They didn't specify.

OpenStudy (anonymous):

ok lets try this 2 beers and a shot cost $10, while 3 beers and 4 shots cost $25 if \(x\) is the price of a beer and \(y\) is the price of a shot, how much to each cost?

OpenStudy (anonymous):

then solve \[2x + y = 10 \]\[3x + 4y = 25\]

OpenStudy (anonymous):

you can solve in many ways. one way is to write the first equation as \[y=10-2x\] and then substitute in to the second equation and get \[3x+4(10-2x)=25\] solve via \[3x+40-8x=25\] \[-5x+40=25\] \[-5x=-15\] \[x=3\]

OpenStudy (anonymous):

so beer cost $3 and therefore shots are $4

OpenStudy (anonymous):

if you think beers and shots will annoy your teacher, change them to hot dogs and hamburgers

OpenStudy (texaschic101):

sounds good to me :) good job @satellite73

OpenStudy (anonymous):

You're a life saver ! Thank you so much! It actually makes sense now too, I've been stuck on this for a while. And yeah, considering I'm only 15, it might haha. Thanks again ! You're amazing !

OpenStudy (anonymous):

yw